21 December 2007

RHIC Startup:  Discussion of the reasons for all the 6 lost fills yesterday.  The first two were due to ripple on the power supplies.  Carl pulled a card to fix after the first fill was lost, but did not complete the repairs until after the second fill was lost.  The third and fourth fill was lost when a procedural error resulted in a sequencer being run with no separation bumps.  The fifth and sixth fills were lost due to an instability produced when the Landau cavity voltage comes on.

Yellow efficiency remains good, but Blue efficiency drops throughout the ramp and transition.  It is not clear why.  The same losses appear in 6 bunch ramps, so it does not depend on the number of bunches. 

There is a smoke alarm in the Blue QLI system.  Yugang would like to reprogram the alarm so that it doesn't take down the beams.  There was a question of whether the alarm time could be lengthened, since it appears to be a transient causing the false alarm, and Ray Karol needs to make this assessment.  The repair would keep beam off for approx 2 hours, so it was put on hold awaiting any unscheduled downtime.  Paul will find out whether the reprogramming can be accomplished over the weekend or holiday, should the downtime occur then.

 There has been a lack of communication between the RF group and the OC regarding the status of the RF.  The RF has been left in CW mode causing unnecessary delays in getting beam going.  This communication needs to be improved, and perhaps have an alarm in the MCR showing the RF status. 

STAR needs to replace a bad HV supply card.  They are waiting for delivery of the new card.  As soon as the new card arrives, then at the end of the subsequent fill they will make a 15 minute access to replace HV cards.

Mike Brennan needs a half hour between stores to work on RF phasing.  This work can probably occur during the STAR access.

The OC worked with PHENIX this morning to study correlations between PHENIX background scintillator S2 and the PHENIX Muon ID currents.  There seems to be fair correlation, so the S2 rate will be used to tune up the PHENIX backgrounds. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020